OpenStudy (anonymous):

On Monday, Cathy read 1/6 of her book, and on Tuesday, she read 2/6 of the book. How much has she read?

OpenStudy (mathmath333):

total amount ,she readed till now would be \[1-\frac{ 1 }{ 6}-\frac{ 2 }{ 6 }\\=\frac{ 6-1-2 }{ 6 }\\=\frac{ 3 }{ 6}\\=\frac{ 1 }{ 2 }\]